# Uncertainty of magnetic field in 3D NLTE inversions

### v2

@@
Uncertainty of magnetic field in 3D NLTE inversions
@@
We describe our approach to solve the problem of ensuring the solenoidality of the magnetic field vector in three-dimensional (3D) inversions, as well as the estimation of the uncertainty in the inferred magnetic field. The solenoidality of the magnetic field vector is often disregarded in the inversion of spectropolarimetric data due to limitations in the traditional one-dimensional inversion techniques. We propose a method to ensure the solenoidal condition in 3D inversions based on our meshfree approach. The increase in dimensionality with respect to the 1D inversion techniques is such that some of the traditional methods to determine the uncertainties become unfeasible. We propose a method based on a Monte Carlo approach to determine the uncertainty of the magnetic field inference. Due to the physics of the problem, we can compute the uncertainty increasing the total required computational time by just a factor of about two. We also propose a metric to quantify the uncertainty to describe the degree of confidence of the magnetic field inference. Finally, we perform a numerical experiment to demonstrate the feasibility of both the method and the metric proposed to quantify the uncertainty.
@@ 1 Introduction
The magnetic field plays a key role in the solar upper atmosphere, driving its dynamics and activity. Spectropolarimetry and inversion techniques are a powerful tool to unravel the properties of these magnetic fields (see, e.g., the review by de la Cruz Rodríguez & van Noort, 2017). However, inferring the 3D distribution of the magnetic field vector from spectropolarimetric observations requires sophisticated methods to account for the intricate interplay of physical effects producing the polarization signals, and thus it remains a significant challenge. Spectral lines in the upper solar atmosphere typically form out of the local thermodynamic equilibrium (NLTE) conditions, and the information carried in their spectropolarimetric profiles, influenced by the Zeeman and Hanle effects, is much richer than that of the photospheric lines forming in, or close to, LTE conditions. On the one hand, the nonlocal and nonlinear coupling introduces considerable challenges for both the physical interpretation and numerical modeling of these spectropolarimetric signals. On the other hand, it offers a framework to infer the magnetic field vector and other quantities throughout the entire 3D geometry in a physically consistent manner.
@@
It has long been recognized that traditional pixel-by-pixel inversion techniques, while providing significant insight on the magnetic field, seem insufficient for the accurate and robust interpretation of spectropolarimetric observations because they fail to account for the spatial coupling of physical quantities across different plasma regions (see, e.g., van Noort, 2012). Several recent studies have incorporated spatial coupling into inversion techniques to address this limitation. Asensio Ramos & de la Cruz Rodríguez (2015) introduced sparsity regularization and showed its potential by implementing it in a Milne-Eddington inversion. de la Cruz Rodríguez (2019) considered the spatial coupling due to the combination of observations with different spatial sampling and instrumental point spread functions and presented their method in a Milne-Eddington inversion. The potential of machine learning techniques such as neural fields has also been explored (Díaz Baso et al., 2024) and implemented in the spatially coupled application of the weak-field approximation of the Zeeman effect.
@@
Although these approaches entail a significant advance toward the more robust inference of the 3D spatial distribution of the magnetic field vector, they still remain limited regarding a possible application to upper-chromosphere spectral lines because they only consider the spatial coupling of thermodynamic quantities while neglecting radiative interactions between matter in different regions in the 3D plasma. Forward-modeling studies, such as those using the PORTA code (Štěpán & Trujillo Bueno, 2013), have demonstrated the importance of including radiative coupling to capture the full complexity of the radiative transfer (RT). Without accounting for these effects, traditional methods may not be able to fully resolve the intricacies of 3D magnetic field topologies, remaining mostly restricted to the estimation of the line-of-sight (LOS) component of the magnetic field in strongly magnetized active regions.
@@
To consider the spatial coupling in full, Štěpán et al. (2022) proposed a new framework for the 3D inversion of spectropolarimetric data, allowing for the inclusion of all relevant physical processes, including RT. The method is formulated as a meshfree and unconstrained minimization problem that includes stochastic ingredients. The unconstrained nature of the problem avoids the need for self-consistency (i.e., the solution of the forward problem at every step of the inversion process), which instead appears in the loss function as a regularization. The method’s stochastic nature helps avoid local minima and saddle points in the global loss function. Additional constraints, such as preservation of the magnetic field solenoidality, ${B}=0$, hydrostatic equilibrium, etc., can be imposed via regularization terms in the loss function. We have recently implemented these ideas in the POLARIS code, which we have used throughout this investigation.
@@
Another critical aspect of the inference of the magnetic field is that of the uncertainties. Not only is the inversion problem ill-posed, but observations have limited spectral, spatial, and temporal resolution and are affected by noise. Without a proper quantification of the uncertainty, the inference is just one of the many possible solutions compatible with the data. A point estimated error based on the response function can be derived from the response functions (Sánchez Almeida, 1997), which provide an indication of how well constrained a node value is relative to the others (Milić & van Noort, 2018). However, while straightforward and relatively computationally cheap in a typical 1D inversion, such an approach is completely impractical in our meshfree 3D method. A much better estimation of the uncertainty can be achieved using statistical techniques, such as Bayesian methods (Asensio Ramos et al., 2007). While it has been shown that neural networks can help with the acceleration of these time-consuming techniques (Díaz Baso et al., 2022), such an approach is impractical in 3D. A different statistical approach consists of performing a significant number of inversions, introducing some randomization in the initial model and/or in the spectropolarimetric profiles (Westendorp Plaza et al., 2001; del Toro Iniesta & Ruiz Cobo, 2016; Sainz Dalda & De Pontieu, 2023), with the uncertainty given by the standard deviation among the whole ensemble of inversions. While a priori this approach also seems impractical for the very time-consuming inversions, we will show how it can be implemented to determine the uncertainty of the magnetic field inference under certain reasonable assumptions.
@@
This paper has two primary goals. First, in Sect. 2, we introduce a method capable to ensure the exact preservation of the solenoidal nature of the magnetic field throughout the entire domain, allowing us to avoid the addition of the solenoidality regularization term in the loss function. This is made possible by the way that our 3D inversion code, POLARIS, treats the model variables. Second, in Sect. 3, we address the challenging task of estimating the uncertainties of the inferred magnetic field vector. This is particularly challenging in 3D models due to the large number of model parameters. We present a simple method for estimating the uncertainties of the magnetic field vector, demonstrating how to derive its probability distribution at any point in the model. Additionally, we propose a quantitative approach for assessing the relative uncertainty of the inversion of the magnetic field across different regions, providing a practical framework for the visualization of the magnetic field in 3D geometry. In Sect. 4, we apply the above-mentioned methods in a numerical experiment in which we invert a known magnetic field in a 3D volume. Finally, our conclusions are summarized in Sect. 5.
@@ 2 Exact preservation of the magnetic field solenoidality
Figure 1: Line-center image of the $5\times 5\times 5\;\mathrm{Mm}^{3}$ model prominence just above the solar limb used in this paper, observed from Earth with a field of view of $20^{\prime\prime}\times 20^{\prime\prime}$, at the line center of the academic line at 5000 Å. The disk illumination is given by the disk center intensity and center-to-limb variation in Allen (1976).
@@
Figure 5: Histogram of the magnetic field components at the position with normalized coordinates ${r}=(0,0,0.5)$ in the 3D domain for 100 different inverted models. The different panels correspond to the magnetic field vector component along the $x$, $y$, and $z$ axes in the model, from top to bottom, respectively. The vertical red line indicates the true value of the magnetic field vector. The relative uncertainty for each component at this position, as defined in Eqs. (2)–(3), are: $U_{x}=0.97$, $U_{y}=0.50$, $U_{z}=0.60$, and $U_{xyz}=0.66$.
Figure 6: Variation of the magnetic field vector components along the LOS. From left to right we show the magnetic field vector component along the $x$, $y$, and $z$ axes in the model, respectively. The different rows correspond to different positions in the FOV whose location in normalized coordinates, $P_{i}=(X,Z)$, are indicated to the left of each row. The blue curve shows the true magnetic field and the red curve the result of the inversion. The dark and light orange-shaded regions indicate the 68 % and 95 % confidence intervals of the inversion, respectively.
@@
The Maxwell equation ${B}=0$ represents a fundamental condition that must be satisfied in any physical model. This condition, known as solenoidality, is naturally imposed in 3D magnetohydrodynamic (MHD) simulations. However, it is typically neglected in NLTE inversion theory because standard 1D inversions do not readily accommodate it. Nevertheless, certain techniques designed to minimize the current density and ${B}$ in the photosphere have been employed in the LTE regime to solve the $180^{\circ}$ ambiguity (e.g., Metcalf et al., 2006). Recently, Borrero et al. (2024) introduced a method to incorporate this constraint by computing one of the magnetic field vector components from the other two to satisfy ${B}=0$.
@@
In the code POLARIS, we implement the ideas presented in Štěpán et al. (2022) using a meshfree approach, eliminating the need for a traditional grid. This design facilitates the representation of a quantity $\theta_{m}$ at a spatial point ${r}\equiv(x,y,z)$ through the expression
@@ 3 Estimation of the error of the inverted magnetic field
Exploring the entire parameter space to identify consistent models that fit the observations would, in principle, allow for the construction of probability distributions for individual parameters and the derived physical quantities. However, in the context of 3D inversions, the vast number of model parameters makes such an exhaustive exploration computationally impossible. In this paper, we address this challenge by narrowing our investigation to the magnetic field, which offers a more tractable problem compared to other variables. This simplification is justified because changes in the magnetic field significantly affect the goodness-of-fit parameter $\chi^{2}$, while typically having a negligible impact on the thermal structure of the inverted model. Specifically, fixing the thermal structure and the $J^{K}_{Q}$ tensors, i.e., our atomic-like variables as defined in the Štěpán et al. (2022) framework, results in minimal changes to the NLTE consistency of the model under variations in the magnetic field. While this does not apply to the $J^{1}_{Q}$ components, which depend exclusively on the circular polarization, their impact on the problem is generally minor. Even in cases in which these tensors impact the solution, the proposed approach remains effective for rough error estimation. By taking advantage of this property, we employ the well-established Monte Carlo (MC) strategy outlined in Sect. 15.6.2 of Press et al. (2007). This approach enables a systematic and computationally efficient estimation of uncertainties specific to the magnetic field. The algorithm can be summarized as:
@@
    Synthesize the FOV spectra from this model, creating synthetic data that will serve as the reference for the uncertainty estimation process.
@@
As a result, we obtain up to $N$ of the ${A}$ vector potential configurations that are consistent with the observations. [^1] Any inverted model from an inversion failing to converge should be excluded from the analysis to ensure the reliability of the results. Each configuration corresponds to a magnetic field ${B}=\nabla\times\@vec{A}$, whose Cartesian components, $B_{i}(\@vec{r})$, can be calculated at any point ${r}$ within the domain. The resulting ensemble of magnetic fields provides a distribution that can be used to estimate the uncertainties as discussed in the following section.
We note that sampling the $c_{ijk}$ coefficients from a normal distribution ensures that the distribution of initial $A_{i}$ components at every point is also sampled from a normal distribution with zero mean.
@@ 3.1 Quantification of uncertainties
The most comprehensive representation of the magnetic field uncertainty at a specific spatial point in the 3D domain is the distribution of the magnetic field vector obtained using the Monte Carlo (MC) method described above. If this distribution is unimodal and well-approximated by a 3D normal distribution, the uncertainty can be effectively characterized by the standard deviation, $\sigma(B_{i}(\@vec{r}))$, of the marginal distributions for each Cartesian component of the field.
Visualizing the inferred magnetic field is clearly more challenging in 3D than in 1D inversions, particularly when incorporating error bars. It is thus useful to define a scalar quantity to quantify the relative reliability of the inversion across different regions in the domain. One such measure is the coefficient of variation, $C_{i}=\sigma(B_{i}(\@vec{r}))/\mu(B_{i}(\@vec{r}))$, where $\mu(B_{i}(\@vec{r}))$ stands for the mean value. The absolute value of this quantity decreases as the certainty of the inversion increases. However, a limitation of this measure is that it can take both positive and negative values and that it can become arbitrarily large. To overcome this issue, we define the relative uncertainty using the standard deviation and the root mean square as
@@ 3.3 Note on ambiguities in 3D problems
Ambiguities in the solar spectropolarimetric inverse problem arise from the inherent degeneracies in the mapping of the magnetic field vector to the observed polarization signals. These ambiguities can appear for both the Zeeman and Hanle effects (see, e.g., Landi Degl’Innocenti & Landolfi, 2004; Asensio Ramos et al., 2008). Both effects depend on the geometry of the magnetic field and the line of sight, and in the case of the Hanle effect also on the illumination conditions. Different magnetic field geometries in specific configurations produce indistinguishable polarization signatures in the emergent radiation, leading to multiple plausible solutions for the same observed data.
Among the primary causes of these ambiguities are symmetries in the radiative transfer problem under simplified assumptions. For example, in the constant-property slab approximations, the radiative transfer equation is solved assuming that local plasma conditions and the cylindrically symmetric illumination dominate the formation of the spectral line (Asensio Ramos et al., 2008). These assumptions neglect the radiative coupling between distinct regions of the slab, which can break the symmetry of the problem. Furthermore, as for relatively optically thin lines it is often assumed that the sensitivity of the emergent polarization is limited to a single or a few particular magnetic field vectors (in case of the so-called multi-component models).
However, when considering an optically thick medium with full 3D NLTE radiative transfer, we expect the symmetries of the problem to be significantly reduced. In this regime, the radiation field is strongly coupled to both the local and non-local thermodynamic and magnetic conditions. NLTE effects account for departures from equilibrium in the population and quantum coherence of atomic levels, which are influenced by the anisotropy and geometry of the radiation field. This coupling introduces a directional dependence to the emergent radiation sensitive to variations in the magnetic field and its surrounding environment. The inclusion of 3D radiative transfer further enhances this sensitivity by capturing the spatial variations in the illumination, breaking the symmetry found in simpler 1D models. In conclusion, 3D NLTE radiative transfer in optically thick media has the potential to resolve or mitigate many of these degeneracies. This approach thus generally provides a more unique mapping of the magnetic field to the observed polarization signals. While it may not completely eliminate all ambiguities, especially in cases with significant noise levels, it offers a promising path toward reducing their prevalence and improving the reliability of solar spectropolarimetric inversions.

### v3

# The uncertainty of magnetic fields in 3D non-local thermodynamic equilibrium inversions

@@
The uncertainty of magnetic fields in 3D non-local thermodynamic equilibrium inversions
@@
We describe our approach to solving the problem of ensuring the solenoidality of the magnetic field vector in 3D inversions, as well as the estimation of the uncertainty in the inferred magnetic field. The solenoidality of the magnetic field vector is often disregarded in the inversion of spectropolarimetric data due to limitations in the traditional 1D inversion techniques. We propose a method for ensuring the solenoidal condition in 3D inversions based on our mesh-free approach. The increase in dimensionality with respect to the 1D inversion techniques is such that some of the traditional methods for determining the uncertainties become unfeasible. We propose a method based on a Monte Carlo approach for determining the uncertainty of the magnetic field inference. Due to the physics of the problem, we can compute the uncertainty while only increasing the total required computational time by a factor of about 2. We also propose a metric for quantifying the uncertainty that thus describes the degree of confidence of the magnetic field inference. Finally, we perform a numerical experiment to demonstrate the feasibility of both the method and the metric proposed to quantify the uncertainty.
@@ 1 Introduction
The magnetic field plays a key role in the solar upper atmosphere, driving its dynamics and activity. Spectropolarimetry and inversion techniques are powerful tools for unraveling the properties of these magnetic fields (see, e.g., the review by de la Cruz Rodríguez & van Noort, 2017). However, inferring the 3D distribution of the magnetic field vector from spectropolarimetric observations requires sophisticated methods that account for the intricate interplay of the physical effects that produce the polarization signals, and thus it remains a significant challenge. Spectral lines in the upper solar atmosphere typically form out of the local thermodynamic equilibrium (NLTE) conditions, and the information carried in their spectropolarimetric profiles, influenced by the Zeeman and Hanle effects, is much richer than that of the photospheric lines forming in, or close to, local thermodynamic equilibrium (LTE) conditions. On the one hand, the nonlocal and nonlinear coupling introduces considerable challenges for both the physical interpretation and numerical modeling of these spectropolarimetric signals. On the other hand, it offers a framework for inferring the magnetic field vector and other quantities throughout the entire 3D geometry in a physically consistent manner.
@@
It has long been recognized that traditional pixel-by-pixel inversion techniques, while providing significant insight into the magnetic field, are insufficient for the accurate and robust interpretation of spectropolarimetric observations because they fail to account for the spatial coupling of physical quantities across different plasma regions (see, e.g., van Noort, 2012). Several recent studies have incorporated spatial coupling into inversion techniques to address this limitation. Asensio Ramos & de la Cruz Rodríguez (2015) introduced sparsity regularization and showed its potential by implementing it in a Milne-Eddington inversion. de la Cruz Rodríguez (2019) considered the spatial coupling due to the combination of observations with different spatial sampling and instrumental point spread functions and presented their method in a Milne-Eddington inversion. The potential of machine learning techniques such as neural fields has also been explored (Díaz Baso et al., 2025), and such techniques have been implemented in the spatially coupled application of the weak-field approximation of the Zeeman effect.
@@
Although these approaches entail a significant advance toward the more robust inference of the 3D spatial distribution of the magnetic field vector, they remain limited. They cannot easily be applied to upper-chromosphere spectral lines because they only consider the spatial coupling of thermodynamic quantities, neglecting radiative interactions between matter in different regions in the 3D plasma. Forward-modeling studies, such as those using the PORTA code (Štěpán & Trujillo Bueno, 2013), have demonstrated the importance of including radiative coupling to capture the full complexity of the radiative transfer (RT). Without accounting for these effects, traditional methods may not be able to fully resolve the intricacies of 3D magnetic field topologies, remaining mostly restricted to the estimation of the line-of-sight (LOS) component of the magnetic field in strongly magnetized active regions.
@@
To consider the spatial coupling in full, Štěpán et al. (2022) proposed a new framework for the 3D inversion of spectropolarimetric data, allowing for the inclusion of all relevant physical processes, including RT. The method is formulated as a mesh-free and unconstrained minimization problem that includes stochastic ingredients. The unconstrained nature of the problem avoids the need for self-consistency (i.e., the solution of the forward problem at every step of the inversion process), which instead appears in the loss function as a regularization. The method’s stochastic nature helps avoid local minima and saddle points in the global loss function. Additional constraints, such as preservation of the magnetic field solenoidality, ${B}=0$, and hydrostatic equilibrium, can be imposed via regularization terms in the loss function. We have recently implemented these ideas in the POLARIS code, which we used throughout this investigation.
@@
Another critical aspect of the inference of the magnetic field is that of the uncertainties. Not only is the inversion problem ill-posed, but observations have limited spectral, spatial, and temporal resolution and are affected by noise. Without a proper quantification of the uncertainty, the inference is just one of the many possible solutions compatible with the data. A point estimated error can be derived from the response functions(Sánchez Almeida, 1997), which provide an indication of how well constrained a node value is relative to the others (Milić & van Noort, 2018). However, while straightforward and relatively computationally cheap in a typical 1D inversion, such an approach is impractical in our mesh-free 3D method. A much better estimation of the uncertainty can be achieved using statistical techniques, such as Bayesian methods (Asensio Ramos et al., 2007). While it has been shown that neural networks can help with the acceleration of these time-consuming techniques (Díaz Baso et al., 2022), such an approach is impractical in 3D. A different statistical approach consists of performing a significant number of inversions and introducing some randomization in the initial model and/or in the spectropolarimetric profiles (Westendorp Plaza et al., 2001; del Toro Iniesta & Ruiz Cobo, 2016; Sainz Dalda & De Pontieu, 2023), with the uncertainty given by the standard deviation of the whole ensemble of inversions. While a priori this approach also seems impractical for the very time-consuming inversions, we show how it can be implemented to determine the uncertainty of the magnetic field inference under certain reasonable assumptions.
@@
In Sect. 2 we introduce a method capable of ensuring the exact preservation of the solenoidal nature of the magnetic field throughout the entire domain, which allows us to avoid the addition of the solenoidality regularization term in the loss function. This is made possible by the way that our 3D inversion code, POLARIS, treats the model variables. In Sect. 3 we address the challenging task of estimating the uncertainties of the inferred magnetic field vector. This is particularly challenging in 3D models due to the large number of model parameters. We present a simple method for estimating the uncertainties of the magnetic field vector, demonstrating how to derive its probability distribution at any point in the model. Additionally, we propose a quantitative approach for assessing the relative uncertainty of the inversion of the magnetic field across different regions, providing a practical framework for the visualization of the magnetic field in 3D geometry. In Sect. 4 we apply the abovementioned methods in a numerical experiment in which we invert a known magnetic field in a 3D volume. Finally, our conclusions are summarized in Sect. 5.
@@ 2 Exact preservation of the magnetic field solenoidality
Figure 1: Line-center image of the $5\times 5\times 5\;\mathrm{Mm}^{3}$ model prominence just above the solar limb used in this paper, observed from Earth with a FOV of $20^{\prime\prime}\times 20^{\prime\prime}$, at the line center of the academic line at 5000 Å. The disk illumination is given by the disk center intensity and center-to-limb variation in Allen (1976).
@@
Figure 5: Histogram of the magnetic field components at the position with normalized coordinates ${r}=(0,0,0.5)$ in the 3D domain for 100 different inverted models. The different panels correspond to the magnetic field vector component along the $x$, $y$, and $z$ axes in the model, from top to bottom, respectively. The vertical red line indicates the true value of the magnetic field vector. The relative uncertainty for each component at this position, as defined in Eqs. (2) and (3), are $U_{x}=0.97$, $U_{y}=0.50$, $U_{z}=0.60$, and $U_{xyz}=0.66$.
Figure 6: Variation in the magnetic field vector components along the LOS. From top to bottom, we show the magnetic field vector component along the $x$, $y$, and $z$ axes in the model, respectively. The different columns correspond to different positions in the FOV, whose location in normalized coordinates, $P_{i}=(X,Z)$, is indicated to the left of each row. The blue curve shows the true magnetic field and the red curve the result of the inversion. The dark and light orange-shaded regions indicate the 68 % and 95 % confidence intervals of the inversion, respectively.
@@
The Maxwell equation ${B}=0$ represents a fundamental condition that must be satisfied in any physical model. This condition, known as solenoidality, is naturally imposed in 3D magnetohydrodynamic (MHD) simulations. However, it is typically neglected in NLTE inversion theory because standard 1D inversions do not readily accommodate it. Nevertheless, certain techniques designed to minimize the current density and ${B}$ in the photosphere have been employed in the LTE regime to solve the $180^{\circ}$ ambiguity (e.g., Metcalf et al., 2006). Recently, Borrero et al. (2024) introduced a method for incorporating this constraint by computing one of the magnetic field vector components from the other two to satisfy ${B}=0$.
@@
In the code POLARIS, we implement the ideas presented in Štěpán et al. (2022) using a mesh-free approach, eliminating the need for a traditional grid. This design facilitates the representation of a quantity $\theta_{m}$ at a spatial point ${r}\equiv(x,y,z)$ through the expression
@@ 3 Estimation of the error of the inverted magnetic field
Exploring the entire parameter space to identify consistent models that fit the observations would, in principle, allow for the construction of probability distributions for individual parameters and the derived physical quantities. However, in the context of 3D inversions, the vast number of model parameters makes such an exhaustive exploration computationally impossible. We address this challenge by narrowing our investigation to the magnetic field, which offers a more tractable problem compared to other variables. This simplification is justified because changes in the magnetic field significantly affect the goodness-of-fit parameter $\chi^{2}$, while typically having a negligible impact on the thermal structure of the inverted model. Specifically, fixing the thermal structure and the $J^{K}_{Q}$ tensors, i.e., our atomic-like variables as defined in the Štěpán et al. (2022) framework, results in minimal changes to the NLTE consistency of the model under variations in the magnetic field. While this does not apply to the $J^{1}_{Q}$ components, which depend exclusively on the circular polarization, their impact on the problem is generally minor. Even in cases in which these tensors impact the solution, the proposed approach remains effective for rough error estimation. By taking advantage of this property, we employed the well-established Monte Carlo (MC) strategy outlined in Sect. 15.6.2 of Press et al. (2007). This approach enables a systematic and computationally efficient estimation of uncertainties specific to the magnetic field. The algorithm can be summarized as:
@@
    Synthesize the field of view (FOV) spectra from this model, creating synthetic data that will serve as the reference for the uncertainty estimation process.
@@
As a result, we obtain up to $N$ of the ${A}$ vector potential configurations that are consistent with the observations. [^1] Any inverted model from an inversion failing to converge should be excluded from the analysis to ensure the reliability of the results. Each configuration corresponds to a magnetic field ${B}=\nabla\times\@vec{A}$, whose Cartesian components, $B_{i}(\@vec{r})$, can be calculated at any point ${r}$ within the domain. The resulting ensemble of magnetic fields provides a distribution that can be used to estimate the uncertainties as discussed in the following section. We note that sampling the $c_{ijk}$ coefficients from a normal distribution ensures that the distribution of initial $A_{i}$ components at every point is also sampled from a normal distribution with zero mean.
@@ 3.1 Quantification of uncertainties
The most comprehensive representation of the magnetic field uncertainty at a specific spatial point in the 3D domain is the distribution of the magnetic field vector obtained using the MC method described above. If this distribution is unimodal and well approximated by a 3D normal distribution, the uncertainty can be effectively characterized by the standard deviation, $\sigma(B_{i}(\@vec{r}))$, of the marginal distributions for each Cartesian component of the field.
Visualizing the inferred magnetic field is clearly more challenging in 3D than in 1D inversions, particularly when incorporating error bars. It is thus useful to define a scalar quantity to quantify the relative reliability of the inversion across different regions in the domain. One such measure is the coefficient of variation, $C_{i}=\sigma(B_{i}(\@vec{r}))/\mu(B_{i}(\@vec{r}))$, where $\mu(B_{i}(\@vec{r}))$ stands for the mean value. The absolute value of this quantity decreases as the certainty of the inversion increases. However, a limitation of this measure is that it can take both positive and negative values and that it can become arbitrarily large. To overcome this issue, we defined the relative uncertainty using the standard deviation and the root mean square as
@@ 3.3 Note on ambiguities in 3D problems
Ambiguities in the solar spectropolarimetric inverse problem arise from the inherent degeneracies in the mapping of the magnetic field vector to the observed polarization signals. These ambiguities can appear for both the Zeeman and Hanle effects (see, e.g., Landi Degl’Innocenti & Landolfi, 2004; Asensio Ramos et al., 2008). Both effects depend on the geometry of the magnetic field and the LOS, and in the case of the Hanle effect also on the illumination conditions. Different magnetic field geometries in specific configurations produce indistinguishable polarization signatures in the emergent radiation, leading to multiple plausible solutions for the same observed data.
Among the primary causes of these ambiguities are symmetries in the RT problem under simplified assumptions. For example, in the constant-property slab approximations, the RT equation is solved assuming that local plasma conditions and the cylindrically symmetric illumination dominate the formation of the spectral line (Asensio Ramos et al., 2008). These assumptions neglect the radiative coupling between distinct regions of the slab, which can break the symmetry of the problem. Furthermore, as for relatively optically thin lines, it is often assumed that the sensitivity of the emergent polarization is limited to a single or a few particular magnetic field vectors (in the case of the so-called multicomponent models).
However, when considering an optically thick medium with full 3D NLTE RT, we expect the symmetries of the problem to be significantly reduced. In this regime, the radiation field is strongly coupled to both the local and nonlocal thermodynamic and magnetic conditions. NLTE effects account for departures from equilibrium in the population and quantum coherence of atomic levels, which are influenced by the anisotropy and geometry of the radiation field. This coupling introduces a directional dependence to the emergent radiation sensitive to variations in the magnetic field and its surrounding environment. The inclusion of 3D RT further enhances this sensitivity by capturing the spatial variations in the illumination, breaking the symmetry found in simpler 1D models. In conclusion, 3D NLTE RT in optically thick media has the potential to resolve or mitigate many of these degeneracies. This approach thus generally provides a more unique mapping of the magnetic field to the observed polarization signals. While it may not completely eliminate all ambiguities, especially in cases with significant noise levels, it offers a promising path toward reducing their prevalence and improving the reliability of solar spectropolarimetric inversions.